OpenStudy (anonymous):

\[\sqrt{16r^7s^5}+7r^2\sqrt{r^3s^5}-6rs \sqrt{r^5s^3}\]

OpenStudy (anonymous):

\[\sqrt[5]{16x ^{13}y^8}\times \sqrt[5]{4xyz ^{10}}\]

OpenStudy (anonymous):

o.k are u here?

OpenStudy (anonymous):

so you can look at the whole thin as being under the root

OpenStudy (anonymous):

one way that i tackle these problems is that i like to put the radical as fractions

OpenStudy (anonymous):

\[\sqrt{a}=a ^{1/2}\]

OpenStudy (anonymous):

in this case the nth root is 5, so it would be \[a^{1/5}\]

OpenStudy (anonymous):

http://www.twiddla.com/538411

OpenStudy (anonymous):

click here, no gimmicks I promise.